find the general solution of the equation 2x(y+z^2)p + y(2y+z^2)q = z^3. and deduce that

yz(z^2+yz-2y) =x^2 is a solution.


1
Expert's answer
2021-04-28T16:26:13-0400
"Solution""2x(y+z^2)\frac{dz}{dx}+y(2y+z^2)\frac{dz}{dy}=z^2""\frac{dx}{2x(y+z^2)}=\frac{dy}{y(2y+z^2)}=\frac{dz}{z^2}"




"y\frac{dx}{2xy^2+2xyz^2}=x\frac{dy}{2xy^2+xyz^2}=xy\frac{dz}{xyz^2}""\frac{ydx−xdy−xydz}{2xy2+2xyz2−2xy2−xyz2−xyz2}=\frac{ydx−xdy−xydz}{0}""ydx\u2212xdy\u2212xydz=0""\frac{dx}{x}−\frac{dy}{y}−dz=0""lnx\u2212lny\u2212z=c""z=ln( \frac{x}{y})+C"

2.)



"yz(z^2+2z\u22122y)=x^2""yz^3+2yz^2\u22122y2z=x^2"

Differentiate respect to x:



"3yz^2p+4yzp\u22122y^2p=2x""py(3z^2+4z\u22122y)=2x"

Differentiate respect to y:



"z^3+3z^2yq+2z^2+4yzq\u22124yz\u22122y^2q=0""qy(3z^2+4z\u22122y)+z^3+2z^2\u22124yz=0"

"\therefore" we have:



"3z^2+4z−2y=\frac{2x}{py}""\frac{2xq}{p}+z^3+2z^2−4yz=0"

Since



"p=\frac{1}{x}""q=\frac{−1}{y}"

then:



"\frac{−2x^2}{y}+z^3+2z^2−4yz=0""\u22122x^2+y(z^3+2z^2\u22124yz)=0""yz(z^2+2z\u22124y)=2x^2"

So:



"\frac{z^2+2z−4y}{z^2+2z−2y}=2""z^2+2z\u22124y=2z^2+4z\u22124y""z^2+2z=0"

So we get that the statement can be proved if


"z=0, or""z=\u22122"